OpenStudy (anonymous):

Can someone make sure I'm doing this right?: An archer releases an arrow from a shoulder height of 1.39 m. When the arrow hits the target 18 m away, it hits point A. When the target is removed, the arrow lands 45 m away. Find the maximum height of the arrow along its parabolic path. The top of the target is 1.5 meters from the ground. I used the points (0,1.39) (18,1.26) (45,0) Where c=1.39 I then got the equations: -.13=324a+18b and -1.39=2025a+45b and solved for a and b a=-.0008765 b=.008555 I then used x=-b/2a and got x=4.89

OpenStudy (anonymous):

This is where I stopped because I wasn't sure what else to do

OpenStudy (perl):

the initial angle is zero?

OpenStudy (anonymous):

What is the initial angle?

OpenStudy (perl):

thats not given

OpenStudy (perl):

but you are given some points on the parabolic trajectory

OpenStudy (anonymous):

Hmm I don't know I think your on the right track though, just not sure where to go from here

OpenStudy (perl):

you have points (0, 1.39) (45,0)

OpenStudy (anonymous):

Yeah and I used them.

OpenStudy (perl):

and you have (18, 1.5 - 16/100 )

OpenStudy (perl):

woops, should be 24 cm

OpenStudy (perl):

( 18, 1.26 )

OpenStudy (perl):

ok you can use TI 84 and curve fit your points

OpenStudy (anonymous):

I don't actually have a graphing calculator

OpenStudy (perl):

y = ax^2 + bx + c a = -8.7654 b = .00856 c = 1.39

OpenStudy (anonymous):

Is that all there is? Or do I have to do more?

OpenStudy (perl):

what you have a system of equations

OpenStudy (perl):

you have points {(x,y) : (0,1.39), (18,1.26), (45,0)} plug them into ax^2 + bx + c = y a(0)^2 + b(0) + c = 1.39 a(18)^2 + b(18) + c = 1.26 a(45)^2 + b(45) + c = 0

OpenStudy (perl):

so the first equation tells us that c = 1.39

OpenStudy (perl):

plug that c value into equation 2 and equation 3 a(18)^2 + b(18) + 1.39 = 1.26 a(45)^2 + b(45) + 1.39 = 0

OpenStudy (perl):

324a + 18b = 1.26 - 1.39 2025a + 45b = -1.39

OpenStudy (perl):

324a + 18b = -0.13 2025a + 45b = -1.39

OpenStudy (perl):

can you solve that system of equations?

OpenStudy (anonymous):

Yup just a second

OpenStudy (anonymous):

a=-.0008765

OpenStudy (anonymous):

b=.008555

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

Still haven't figured this out

OpenStudy (perl):

your arithmetic is wrong somewhere

OpenStudy (anonymous):

Well I'm not sure what I'm doing wrong then because I keep getting the same answer

OpenStudy (perl):

we want to solve 324a + 18b = -0.13 2025a + 45b = -1.39 agreed?

OpenStudy (anonymous):

Yup

OpenStudy (perl):

how would you like to solve it, by substitution or elimination?

OpenStudy (anonymous):

I used elimination

OpenStudy (perl):

ok

OpenStudy (perl):

elimination is a bit of work

jimthompson5910 (jim_thompson5910):

I'm getting a=-0.0008765 b=0.008555 c = 1.39 just as you are. Geogebra confirms it

OpenStudy (anonymous):

So then thats right?

OpenStudy (perl):

you used the points (0,1.39) (18,1.26) (45,0) ?

OpenStudy (anonymous):

Yeah

OpenStudy (perl):

oh , i left out x 10 ^-4

OpenStudy (perl):

y = ax^2 + bx + c a = -8.7654 * 10^-4 b = .00856 c = 1.39

OpenStudy (anonymous):

So what would be my next step? I'd have to use x=-b/2a right

OpenStudy (perl):

correct

jimthompson5910 (jim_thompson5910):

the calculation for x = -b/(2a) is off by a digit though x = -b/(2a) x = -(0.008555/(2*(-0.0008765)) x = 4.880205 x = 4.88 if you round to 2 decimal places

OpenStudy (perl):

x = -.008556 / ( 2 * -.0008765432)

OpenStudy (anonymous):

Oh okay. This is the part I got stuck at, how do I find y?

jimthompson5910 (jim_thompson5910):

once you have the x coordinate of the vertex, you plug it into the function that represents the arc to get the y coordinate of the vertex

jimthompson5910 (jim_thompson5910):

so plug it into f(x)

OpenStudy (anonymous):

Just f(x)? So f(4.88)?

jimthompson5910 (jim_thompson5910):

what is f(4.88) equal to?

jimthompson5910 (jim_thompson5910):

remember that f(x) is the quadratic that goes through all 3 points

OpenStudy (anonymous):

I don't know how to find that

jimthompson5910 (jim_thompson5910):

what is f(x) ?

jimthompson5910 (jim_thompson5910):

what is the actual function or quadratic

jimthompson5910 (jim_thompson5910):

you found the values of a,b,c they will be used in f(x) = ax^2 + bx + c

OpenStudy (anonymous):

So then f(4.88)=-.008765x^2+.008555+1.39

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

you forgot about the x attached to b

OpenStudy (anonymous):

oh oops

jimthompson5910 (jim_thompson5910):

and you would replace EVERY x with the 4.88

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Is it 1.22302?

jimthompson5910 (jim_thompson5910):

it's a little short

jimthompson5910 (jim_thompson5910):

I'm getting a slightly larger answer.

OpenStudy (anonymous):

Wait I think I did that wrong. 1.33577947?

jimthompson5910 (jim_thompson5910):

you should have -0.0008765(4.88)^2+0.008555(4.88)+1.39 = 1.41 you had 2 zeros in the first coefficient when it should have been 3

jimthompson5910 (jim_thompson5910):

btw I remember doing this a while back with another person and I made up this drawing

jimthompson5910 (jim_thompson5910):

so that gives a visual summary in a way

OpenStudy (anonymous):

Okay so I got 1.408767606?

OpenStudy (anonymous):

Oh well that makes it a bit easier lol thank you

jimthompson5910 (jim_thompson5910):

that's closer, so if you round what you got, you would have 1.41

OpenStudy (anonymous):

So then the maximum height is (4.88,1.41)

jimthompson5910 (jim_thompson5910):

that is the point where the max height happens the actual max height is 1.41 meters. This is the y coordinate of the vertex point

jimthompson5910 (jim_thompson5910):

the 4.88 is the x coordinate that tells you how far away horizontally the arrow is from the archer

OpenStudy (anonymous):

Oh alright

OpenStudy (anonymous):

And thats all then?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Thank you so much for sticking around and walking me through it!!

jimthompson5910 (jim_thompson5910):

sure thing